OpenStudy (kaylala):

(3^2n times a^3)^2 = ?

OpenStudy (anonymous):

hmm does it have choices? my answer is 9a^6n but i could be wrong

OpenStudy (kaylala):

\[(3^{4n}a ^{3})^{2^{}} = ???\]

OpenStudy (anonymous):

ok sorry i was wrong

OpenStudy (kaylala):

@TjRoDz you're wrong...

OpenStudy (anonymous):

uhh 81a^6n

OpenStudy (kaylala):

still wrong @TjRoDz

OpenStudy (anonymous):

hhhaha my pea sized brain

OpenStudy (kaylala):

help me please @thomaster

thomaster (thomaster):

uhm i think \(\sf\Large a^6\large3^{\Large8n}\)

thomaster (thomaster):

\(\Large (a^n)^m=a^{n*m}\)

OpenStudy (kaylala):

wont it be instead of 3 it would be squared and be 9?

OpenStudy (kaylala):

@thomaster wont it be: instead of 3; it would be squared and be 9?

OpenStudy (kaylala):

i'm confused

thomaster (thomaster):

No \(\Large(3^{4n}a ^3)^2\) 4*2=8 and 3*2=6 \(\Large3^{8n}a ^6~\to~a^63^{8n}\)

OpenStudy (phi):

you can memorize the rule \[ (x^a)^b = x^{ab} \] which means multiply the exponents but you can use this idea: \[ x^2 \text{ means } x \cdot x \] so \[ (3^{4n}a ^{3})^{2} \text{ means } (3^{4n}a ^{3})(3^{4n}a ^{3}) \]

OpenStudy (phi):

you can re-arrange that to \[ (3^{4n}\cdot 3^{4n}\cdot a ^{3}\cdot a ^{3})\]

OpenStudy (phi):

if you don't know about "adding exponents" you could figure out \[ a^3 \cdot a^3 \] by knowing that \(a^3 = a\cdot a \cdot a \) \[ a^3 \cdot a^3 = a\cdot a \cdot a \cdot a\cdot a \cdot a= a^6\]

OpenStudy (phi):

but to do \[ 3^{4n} \cdot 3^{4n} \] you need to know the rule: if you have the same base, add the exponents \[ 3^{4n+4n} \] or \[3^{8n} \]